OpenStudy (anonymous):

Write a polynomial function of minimum degree with real coefficients whose zeros include those listed. Write in standard form. 4, -8, and 2 + 5i I got up to (x-4)(x+8)(2+5i)(2-5i) What do I do next?

OpenStudy (anonymous):

there is a mistake there

OpenStudy (anonymous):

the last to factors should be \[(x-(2+5i))(x-(2+5i))\] there is an easy way to find this however, so it is not as bad as it looks

OpenStudy (anonymous):

*last TWO factors

OpenStudy (anonymous):

\[(x-4)(x+8)\] is routine, what you need is the polynomial of degree two with the two zeros \(2+5i\) and \(2-5i\) right?

OpenStudy (anonymous):

i can show you how to do this more or less in your head if you like

OpenStudy (anonymous):

Oh please do, but do I multiply them all together?

OpenStudy (anonymous):

oh no i wouldn't do that

OpenStudy (anonymous):

first of all we can work backwards that is one way put \[x=2+5i\] subtract 2 and get \[x-2=5i\] square and get \[x^2-4x+4=-25\] add 25 and get \[x^2-4x+29\] as your needed polynomial

OpenStudy (anonymous):

an even easier way to do it, is to memorize a simple fact: if \(a+bi\) is a zero of a quadratic, then it is \(x^2-2ax+a^2+b^2\)

OpenStudy (anonymous):

so if \(2+5i\) is a zero, the quadratic is \(x^2-2\times 2x+2^2+5^2=x^2-4x+29\)

OpenStudy (anonymous):

your math teacher might want you to multiply all that out, and it is not that hard if you do it correctly, group terms etc you will get the same answer of course, but who needs all that aggravation?

OpenStudy (anonymous):

you are not finished the problem, you still have to multiply \[(x-4)(x+8)(x^2-4x+29)\]

OpenStudy (anonymous):

Is there any easy way to multiply that all out?

OpenStudy (anonymous):

you get \[x^4-19 x^2+244 x-928\] with the click of a button other than that, no just have to grind it til you find it

OpenStudy (anonymous):

Thank you soo much :)

OpenStudy (anonymous):

yw